Do each of a-d, both geometrically (you need not be precise) and using calculus. There are only two goods; x is the quantity of one good and y of the other. Your income is I and u(x,y) = xy + x + y.

(a) Px = $2; Py = $1; I = $15. Suppose Py rises to $2. By how much must I increase in order that you be as well off as before?


1
Expert's answer
2021-12-06T16:27:26-0500

Assume the utility function to be:

"U=xy+x+y"

Determining the slope indifference curve:

"dU=xdy+ydx+dx+dy"

Along any indifference curve, "dU=0"

"xdy+ydx+dx+dy=0\\dy(x+1)+dx(y+1)=0\\dy(x+1)=-dx(y+1)\\\frac{dy}{dx}=-\frac{y+1}{x+1}\\MRS_{xy}=-\frac{y+1}{x+1}"

The slope of indifference curve is negative

Differentiating it again

"\frac{d(MRS_{xy})}{dx}=-[\frac{(x+1)\frac{dy}{dx}-(y+1)}{(x+1)^2}]\\\frac{d(MRS_{xy})}{dx}=-[-\frac{(x+1)\frac{y+1}{x+1}-(y+1)}{(x+1)^2}]>0"

Therefore, the indifference curve is strict convex

The budget equation is:

"P_xx+P_yy=I"

Solving this in general:

max: "U=xy+x+y"

subject to: "P_xx+P_yy=I"

The Lagrangian expression is:

"L=xy+x+y+\lambda[I-xP_x-yP_y]"

The first order conditions are:

"\frac{\partial L}{\partial x}=y+1-\lambda P_x=0-----(1)\\\frac{\partial L}{\partial y}=x+1-\lambda P_y=0-----(2)
\\\frac{\partial L}{\partial \lambda}=I=xP_x+yP_y-----(3)"

From (1) and (2):

"\frac{y+1}{x+1}=\frac{P_x}{P_y}\\P_xx+P_x=yP_y+P_y\\P_xx=yP_y+P_y-P_x"

Substitute this in (3):

"yP_y+P_y=P_x+yP_y=I\\2yP_y+P_y-P_x=I\\y=\frac{I+P_x-P_y}{2P_y}"

This is the demand function of good y

"P_yy=\frac{I+P_x-P_y}{2}\\P_xx=yP_y+P_y-P_x\\P_xx=\frac{I+P_x-P_y}{2}+P_y-P_x\\P_xx=\frac{I+P_x-P_y+2P_y-2P_x}{2}\\P_xx=\frac{I+P_y-P_x}{2}"

"x=\frac{I+P_y-P_x}{2P_x}"

This is the demand function for good x


The prices are $2 and $1 for good x and good y respectively and the income is $15

Thus, the budget function is:

"2x+y=15"

max: "U=xy+x+y"

subject to: "2x+y=15"


"x=\frac{I+P_y-P_x}{2P_x}=\frac{15+1-2}{2*2}=3.5\\y=\frac{I+P_x-P_y}{2P_y}=\frac{15+2-1}{2*1}=8"

The demand function for x is 3.5 and y is 8 units

Now, price of y rises to $2 but individual wants to consume initial level of utility

The new budget equation is:

"2x+2y=I'"

Where "I'" is unknown

The initial level of utility is:

"U=3.5*8+3.5+8=39.5\approx40"

Now, "x^s=\frac{I'+2-2}{2*2}=\frac{I'}{4}"

and

"y^s=\frac{I'+2-2}{2*2}=\frac{I'}{4}"

"U=\frac{I'^2}{16}+\frac{I'}{4}+\frac{I'}{4}=40"

Then:

"\frac{I'^2+8I'}{16}=40\\I'^2+8I'=640\\I'^2+8I'-640=0"

You can solve this by the quadratic formula:

"I'=\frac{-8\pm\sqrt{64+(4*640)}}{2}=\frac{-8\pm51.22}{2}"

Income should not be negative, so "I'=21.61"

Therefore, income should be increased by:

"21.61-15=6.61"